OpenStudy (anonymous):

Write the equation of the line that passes through the two points. (-2,6), (3,0)

OpenStudy (saifoo.khan):

Slope?

OpenStudy (anonymous):

Umm...

OpenStudy (saifoo.khan):

let him Solve too.

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

he is so slow saif

OpenStudy (saifoo.khan):

tht isnt the matter man! he will learn this way!

OpenStudy (anonymous):

(y2 - y1)/(x2 - x1) = m = slope

OpenStudy (anonymous):

yes i did that thanks for the reminder

OpenStudy (anonymous):

so i got -6/5 for the slope

OpenStudy (anonymous):

now use point slope formula y - y1 = m(x-x1) with either points (x1,y1) that were given.

OpenStudy (anonymous):

since you know m and a point, then you'll have an equation so solve and that's the equation...

OpenStudy (anonymous):

thanks.